OpenStudy (anonymous):

Solve: -10=3y-5(y-2) 10 4 -10 -4

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

-10=3y-5y+10 -10-10=3y-5y now solve for y...can u do it?

OpenStudy (anonymous):

have u undrstood how have i done?....the brackt term will be multiplied by 5

OpenStudy (anonymous):

-20=-2y minus gets cancelled, y=20/2 y=10

OpenStudy (anonymous):

Yes I understood that thank you :)

OpenStudy (petiteme):

-10=3y-5(y-2) distribute the -5 -10=3y-5y+10 Add like terms -10=-2y +10 transpose 10 -10 -10 = -2y -20=-2y Divide both sides by -2 to get y y=10 :)